OpenStudy (anonymous):

Solve for m. m – (–24) = –17 A. 7 B. 41 C. –7 D. –41

OpenStudy (alexandervonhumboldt2):

m+24=-17

OpenStudy (alexandervonhumboldt2):

m=-17-24

OpenStudy (alexandervonhumboldt2):

m=-41

OpenStudy (anonymous):

r u shere

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

r u really shere

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

help peeps

Nnesha (nnesha):

okay first you have to distriubte

OpenStudy (alexandervonhumboldt2):

m+24=-17 m=-41

Nnesha (nnesha):

distribute**

Nnesha (nnesha):

 negative on times -24 = ???